OpenStudy (anonymous):

In the following equation, identify the x-intercepts in the graph y = 3x2 + 19x – 40

OpenStudy (lynfran):

just to be clear is this it.. \[3x ^{2}+19x-40\]

OpenStudy (anonymous):

yes

OpenStudy (lynfran):

so we have (x+8)(3x-5) equate each bracket to 0 and solve for x x+8=0 3x-5=0

OpenStudy (lynfran):

what did you get?

OpenStudy (lynfran):

x=-8 and x=5/3

OpenStudy (anonymous):

Thank you so much

OpenStudy (lynfran):

welcome